OpenStudy (anonymous):

What are the solutions to the equation x2 + 4 = 148? A. –11 and 11 B. –2 and 2 C. –12 and 12 D. –4 and 4

OpenStudy (anonymous):

@TwiztTiez @FEARLESS_JOCEY

OpenStudy (anonymous):

i dont know for sure umm @dan815 can you help?

OpenStudy (anonymous):

@Chiko_1278

OpenStudy (anonymous):

do opposite opperation

OpenStudy (anonymous):

what

OpenStudy (anonymous):

and remember what you do on one side do it on the other side

OpenStudy (anonymous):

it dosnt work like that

OpenStudy (anonymous):

or what you try to find?

OpenStudy (anonymous):

the solution which its 72

OpenStudy (anonymous):

wait....ohhh ok gimme a sec..

OpenStudy (anonymous):

i go for B

OpenStudy (anonymous):

have you worked it out yet

OpenStudy (anonymous):

First step is to rearrange the equation, getting everything on one side x^2 - 144 = 0 At this point it's pretty clear what to do Difference of Perfect Squares (x+12)(x-12) = 0 (x+12) = 0, (x-12) = 0 x = -12, 12

OpenStudy (anonymous):

i mean C ....my bad

OpenStudy (anonymous):

yes...no..????...idk????

OpenStudy (anonymous):

did i help?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok tag me whenever you need help..have a good day

OpenStudy (anonymous):

☺☺☺☺☺☺☺